OpenStudy (countrygirl1431):

A cylinder has a height of 5.3 cm and a diameter of 3.0 cm. What is its volume? (Hint: Volume of a cylinder = r2h ) 25.0 cm3 66.2 cm3 149.9 cm3 37.5 cm3

undeadknight26 (undeadknight26):

The formula for volume is actually pir^2h.

OpenStudy (countrygirl1431):

i know i just couldnt put pi

OpenStudy (countrygirl1431):

but r is the radius so couldnt i just do the diameter

undeadknight26 (undeadknight26):

I see...Well the hardest/trickiest thing about this question is: V = pi * r^2 * h... Would it be: V = pi * 3^2 * 5.3 or: V = pi * 1.5^2 * 5.3? Diameter is from one side of the circle to the other while radius is from one side to the middle (Half the diameter).

OpenStudy (countrygirl1431):

right so the diameter =r^2

undeadknight26 (undeadknight26):

Mhmm

OpenStudy (countrygirl1431):

so would have to square the diameter

OpenStudy (countrygirl1431):

so it would be 149.9 right?

undeadknight26 (undeadknight26):

I don't think so...

OpenStudy (countrygirl1431):

thats what i got

undeadknight26 (undeadknight26):

Did you do 1.5 x 2? or 1.5 x 1.5? 1.5 x 2 = 3 1.5 x 1.5 = 2.25

OpenStudy (countrygirl1431):

37.5

undeadknight26 (undeadknight26):

:D I would believe that is your answer :D

OpenStudy (countrygirl1431):

ok thanks

undeadknight26 (undeadknight26):

Any time m8